OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. Answer -13 4 13 -12

OpenStudy (johnweldon1993):

Well lets see...we use \[\frac{ y_2 - y_1 }{ x_2 - x_1 }\] to find slope...we know the numerator of the slope = -3...so lets find out what we have y2 - y1 = -3 - 6 = -9 how do we get from -9 to -3? we divide by 3....so this mean....to find the denominator...we need to find what number when subtracted from 10......and divided by 3....= 2 x2 - x1 = 10 - x we know \[\frac{ 10 - x }{ 3 } = 2\] so multiply both sides by 3 \[10 - x = 6\] subtract 10 from both sides \[-x = -4\] and multiply by -1 x = 4 so the 'x' must equal 4

OpenStudy (johnweldon1993):

Now lets check it \[\frac{ y_2 - y_1 }{ x_2 - x_1 }\] \[\frac{ -3 - 6 }{ 10 - 4 }\] \[\frac{ -9 }{ 6 }\] simplified down \[-\frac{ 3 }{ 2 }\] So it works out